OpenStudy (anonymous):

equation:x^(2)y'-2xy=cos(1/x) I have to solve it using legrand's variation method or integral factor method.Could someone help me,please?

OpenStudy (anonymous):

First arrange it in standard form so y' stands alone /:x^2 y'-2y/x=cos(1/x) / x^2

OpenStudy (anonymous):

Than the integrating factor is exp(integral -2/x)=exp(-2lnx)=x^-2

OpenStudy (anonymous):

Multiplying with it will give you y'x^-2-2yx^-3=cos(1/x)x^-4

OpenStudy (anonymous):

Here you notice that y'x^-2-2yx^-3=(yx^-2)'

OpenStudy (anonymous):

So (yx^-2)'=cos(1/x)x^-4 pr yx^-2=integral cos(1/x)x^-4

OpenStudy (anonymous):

Now I have to take out my pizza from the oven .-)

OpenStudy (anonymous):

now Im stuck a bit. We need to find the integral cos(1/x)x^-4

OpenStudy (anonymous):

Sorry I have no idea at all how to do that.

OpenStudy (anonymous):

ok, i m starting to look your soultion

OpenStudy (anonymous):

Here comes the cavalry! :-)

OpenStudy (anonymous):

:) but you re saying that x^(-2) is the integral factor ?

OpenStudy (anonymous):

Yes. I dont know legrand's method. Maybe amistre is doing that right now

OpenStudy (anonymous):

hmm,you re right it is the integral factor,and we have to integrate cos(1/x)/x^4,which is imposible to inte grate with elementary functions :(

OpenStudy (amistre64):

\[x^2y'-2xy=cos(1/x)\] \[\frac{1}{x^2}(x^2y'-2xy=cos(1/x))\] \[y'-\frac{2}{x}y=\frac{cos(1/x)}{x^2}\] \[exp(-2ln|x|)\ (y'-\frac{2}{x}y=\frac{cos(1/x)}{x^2})\] \[exp(-2ln|x|) y'+\ \frac{2}{x}exp(-2ln|x|)\ y=exp(-2ln|x|)\frac{cos(1/x)}{x^2}\] \[exp(-2ln|x|) y\ =\ \int exp(-2ln|x|)\frac{cos(1/x)}{x^2}\ dx\] \[y\ =exp(2ln|x|) \int exp(-2ln|x|)\frac{cos(1/x)}{x^2}\ dx+C_1exp(-2ln|x|)\] \[y\ =x^2 \int \frac{cos(1/x)}{x^4}\ dx+C_1x^{-2}\]

OpenStudy (amistre64):

x^4 can be partial fractioned i believe

OpenStudy (amistre64):

\[\frac{cos(1/x)}{x^4}=\frac{a}{x}+\frac{b}{x^2}+\frac{c}{x^3}+\frac{d}{x^4}\] \[cos(1/x)=ax^3+bx^2+cx+d\] maybe or some other trick

OpenStudy (amistre64):

integrate by parts perhaps

OpenStudy (anonymous):

ou,yes , i had forgotten of partial fractions

OpenStudy (amistre64):

\[\int\frac{1}{x^2}\frac{cos(\frac{1}{x})}{x^2}dx\] \[=-\frac{1}{x^2}sin(\frac{1}{x})-\int \frac{2}{x^3}sin(\frac{1}{x})dx\] \[=-\frac{1}{x^2}sin(\frac{1}{x})-2\int \frac{1}{x}\frac{sin(\frac{1}{x})}{x^2}dx\] \[=-\frac{1}{x^2}sin(\frac{1}{x})-2(-\frac{1}{x}cos(\frac{1}{x})+\int \frac{}{})\] oy this gives me a head ache ... its gonna be a mess of an integration by parts ...

OpenStudy (amistre64):

i dont know the other method either ...

OpenStudy (anonymous):

ok,thanks,the other method in the end has the same integration problem.